# COREval: A Comprehensive and Objective Benchmark for Evaluating the Remote Sensing Capabilities of Large Vision-Language Models

### v1

@@ 3.3 Evaluation Strategy
       Model ILC SII CID AttR AssR CSR
       General-domain Large Vision-Language Models
Qwen2-VL-7B[41] 0.830 0.682 0.703 0.243 0.233 0.910
InternVL2-8B[4] 0.806 0.625 0.666 0.480 0.278 0.822
LLaVA-1.6-7B[23] 0.746 0.543 0.630 0.230 0.155 0.718
LLama3.2-11B[9] 0.749 0.524 0.614 0.290 0.158 0.808
GLM-4V-9B[10] 0.788 0.572 0.655 0.187 0.010 0.846
DeepSeek-VL-7B[26] 0.796 0.638 0.628 0.270 0.060 0.860
MiniCPM-V-2.5[46] 0.785 0.603 0.653 0.247 0.208 0.882
Phi3-Vision[1] 0.750 0.547 0.584 0.280 0.075 0.714
Remote Sensing Large Vision-Language Models
GeoChat[15] 0.726 0.508 0.251 0.327 0.138 0.696
LHRS-Bot[29] 0.708 0.317 0.181 0.267 0.230 0.574
LHRS-Bot-nova[29] 0.768 0.526 0.262 0.327 0.143 0.578
RN50 0.717 0.328 0.504 0.307 0.225 0.416
ViT-B 0.753 0.348 0.338 0.287 0.185 0.500
RemoteCLIP[22] ViT-L 0.709 0.355 0.514 0.293 0.140 0.736
ViT-B 0.757 0.311 0.541 0.327 0.243 0.828
ViT-B_RET-2 0.772 0.255 0.447 0.337 0.188 0.766
ViT-L 0.749 0.299 0.571 0.327 0.230 0.904
ViT-L-336 0.762 0.333 0.598 0.347 0.170 0.906
GeoRSCLIP[49] ViT-H 0.763 0.331 0.404 0.370 0.285 0.928
Table 2: Evaluation results for L-2 dimensions. Abbreviations adopted: ILC for Image-level Comprehension; SII for Single-instance Identification; CID for Cross-instance Discernment; AttR for Attribute Reasoning; AssR for Asessment Reasoning; CSR for common sense Reasoning. The best (second best) is in bold (underline).
@@ 4 Evaluation Results
  Image-Level Comprehension Single-Instance Identification Cross-Instance Discernment
@@
Qwen2-VL-7B[41] 0.720 0.527 0.796 0.935 0.975 0.970 0.510 0.775 0.910 0.965 0.310 0.560 0.790 0.490 0.780
InternVL2-8B[4] 0.755 0.487 0.452 0.936 0.955 0.770 0.390 0.640 0.900 0.979 0.065 0.800 0.815 0.401 0.720
LLaVA-1.6-7B[23] 0.455 0.410 0.401 0.903 0.920 0.720 0.400 0.635 0.730 0.951 0.240 0.330 0.771 0.401 0.665
LLama3.2-11B[9] 0.555 0.356 0.624 0.901 0.890 0.890 0.380 0.700 0.900 0.896 0.005 0.185 0.790 0.318 0.665
GLM-4V-9B[10] 0.545 0.434 0.675 0.930 0.935 0.970 0.470 0.440 0.920 0.965 0.000 0.550 0.854 0.344 0.695
DeepSeek-VL-7B[26] 0.530 0.469 0.669 0.924 0.895 0.920 0.330 0.770 0.950 0.972 0.110 0.555 0.756 0.369 0.700
MiniCPM-V-2.5[46] 0.650 0.396 0.675 0.925 0.970 0.900 0.400 0.630 0.905 0.958 0.020 0.655 0.776 0.408 0.720
Phi3-Vision[1] 0.530 0.439 0.497 0.886 0.900 0.650 0.380 0.770 0.880 0.910 0.055 0.340 0.717 0.414 0.580
@@
GeoChat[15] 0.335 0.367 0.439 0.902 0.830 0.790 0.295 0.775 0.820 0.708 0.260 0.105 - 0.318 0.455
LHRS-Bot[29] 0.300 0.331 0.306 0.920 0.630 0.480 0.290 0.245 0.870 0.347 - 0.075 - 0.178 0.370
LHRS-Bot-nova[29] 0.470 0.406 0.318 0.954 0.830 0.710 0.290 0.660 0.840 0.743 0.285 0.305 - 0.261 0.530
RN50 0.300 0.337 0.306 0.925 0.705 0.490 0.460 0.435 0.460 0.382 - 0.165 0.712 0.268 0.475
ViT-B 0.585 0.337 0.274 0.953 0.765 0.730 0.305 0.350 0.370 0.826 - 0.180 0.380 0.248 0.365
RemoteCLIP[22] ViT-L 0.510 0.303 0.369 0.891 0.775 0.700 0.435 0.295 0.715 0.500 - 0.050 0.688 0.248 0.545
ViT-B 0.660 0.317 0.363 0.945 0.830 0.940 0.210 0.180 0.605 0.535 - 0.085 0.654 0.274 0.635
ViT-B_RET-2 0.645 0.317 0.376 0.967 0.855 0.860 0.205 0.140 0.360 0.424 - 0.145 0.478 0.261 0.560
ViT-L 0.625 0.280 0.446 0.941 0.835 0.930 0.230 0.310 0.570 0.292 - 0.075 0.707 0.325 0.625
ViT-L-336 0.685 0.319 0.427 0.944 0.835 0.940 0.290 0.310 0.775 0.229 - 0.060 0.732 0.274 0.715
GeoRSCLIP[49] ViT-H 0.705 0.303 0.471 0.944 0.860 0.980 0.230 0.360 0.750 0.250 - 0.050 0.361 0.248 0.570
Table 3: Fine-grained evaluation results for Perception. Abbreviations adopted: IM for Image Modality; IQ for Image Quality; MR for Map Recognition; SC for Scene Classification; IC for Image Caption; LR for Landmark Recognitionl; OC for Object Counting; OL for Object Localization; OP for Object Presence; AR for Atrribute Recognition; VG for Visual Grounding; HD for Hallucination Detection; AC for Atrribute Comparison; SR for Spatial Relationship; CD for Change Detection. The best (second best) is in bold (underline).
4.1 Experimental Setup
We evaluated 13 mainstream open-source VLMs based on COREval, which can be categorized into two groups: 8 general-domain VLMs, including Qwen2-VL [41], InternVL2 [4], LLaVA [23], Llama3.2 [9], GLM-4V [10], DeepSeek-VL [26], MiniCPM-V [46], and Phi3-Vision [1], and 5 RSVLMs: GeoChat [15], LHRS-Bot [29], LHRS-Bot-nova [29], RemoteCLIP [22], and GeoRSCLIP [49]. All VLMs in our evaluation employ the same zero-shot prompt strategy. Since most of our problems are MCQs, this format ensures the objectivity and rigor of the evaluation. For Visual Grounding, we set mIoU metric to ensure the accuracy and objectivity of the assessment results.
4.2 Main Results
The overall evaluation results of COREval are depicted in LABEL:fig:eval_dim_results and Tab. 2. Currently, general-domain VLMs, benefiting from extensive training corpora, exhibit a broader range of capabilities, even outperforming RSVLMs in certain abilities pertinent to remote sensing. However, none of the VLMs achieve consistently high performance across the 6 capability dimensions of COREval, highlighting substantial room for improvement in tackling complex remote sensing tasks.
ILC and CSR Dimensions. For the L-2 dimensions of COREval, the results indicate that the majority of general-domain VLMs and RSVLMs exhibit strong capabilities in ILC and CSR, with task accuracy rates in these dimensions reaching approximately up to 80% or higher. This outcome demonstrates that these VLMs generally possess the foundational and critical capabilities for reasoning about common sense knowledge in remote sensing scenarios and comprehending RSIs at the image level.
SII and CID Dimensions. However, for the perception of finer-grained remote sensing targets, the results in Single-instance Identification and Cross-instance Discernment show that the accuracy of most general VLMs falls below 70%, while that of RSVLMs is under 60%, indicating that current VLMs exhibit limitations in perception capabilities towards more complex remote sensing scenarios and targets, necessitating further enhancements.
AttR and AssR Dimensions. Additionally, capabilities in AttR and AssR, which involve the analysis and reasoning of advanced attributes and features relevant to remote sensing, are largely underdeveloped in most existing VLMs. The evaluated models exhibit an accuracy of nearly below 50% in AttR, with AssR performing even worse, falling beneath 30%. However, these capabilities are of significant importance and value for the practical applications of large models in authentic remote sensing scenarios, underscoring the need for greater focus in this area.
Attribute
Reasoning
Assessment
Reasoning
Common Sense
Reasoning
Model TP PP EA RA DD GD SI
       General Large Vision-Language Models
Qwen2-VL-7B[41] 0.350 0.030 0.305 0.160 0.880 0.955 0.880
InternVL2-8B[4] 0.365 0.710 0.185 0.370 0.870 0.790 0.790
LLaVA-1.6-7B[23] 0.275 0.140 0.270 0.040 0.730 0.670 0.790
LLama3.2-11B[9] 0.365 0.140 0.150 0.165 0.790 0.860 0.740
GLM-4V-9B[10] 0.255 0.050 0.010 0.010 0.895 0.830 0.780
DeepSeek-VL-7B[26] 0.315 0.180 0.060 0.060 0.835 0.885 0.860
MiniCPM-V-2.5[46] 0.330 0.080 0.325 0.090 0.900 0.890 0.830
Phi3-Vision[1] 0.350 0.140 0.080 0.070 0.870 0.570 0.690
Remote Sensing Large Vision-Language Models
GeoChat[15] 0.335 0.310 0.115 0.160 0.630 0.805 0.610
LHRS-Bot[29] 0.345 0.110 0.360 0.100 0.555 0.465 0.830
LHRS-Bot-nova[29] 0.325 0.330 0.105 0.180 0.505 0.515 0.850
RN50 0.300 0.320 0.180 0.270 0.250 0.420 0.740
ViT-B 0.310 0.240 0.130 0.240 0.280 0.620 0.700
RemoteCLIP[22] ViT-L 0.335 0.210 0.035 0.245 0.820 0.650 0.740
ViT-B 0.400 0.180 0.175 0.310 0.855 0.845 0.740
ViT-B_RET-2 0.370 0.270 0.145 0.230 0.785 0.795 0.670
ViT-L 0.375 0.230 0.160 0.300 0.970 0.890 0.800
ViT-L-336 0.430 0.180 0.120 0.220 0.945 0.885 0.870
GeoRSCLIP[49] ViT-H 0.400 0.310 0.310 0.260 0.945 0.935 0.880
4.3 Fine-grained Analysis
@@
Tab. 3 and Tab. 4 show the experimental results of general-domain VLMs and RSVLMs across all L-3 leaf tasks on COREval.
@@ 4.3.1 Model Dimensions
General-domain VLMs outperform RSVLMs in perception tasks. General-domain VLMs demonstrate superior performances in most tasks compared to RSVLMs. Notably, Qwen2-VL achieves the highest accuracy in a majority of remote sensing tasks, exceeding 75% in MR, IC, OL, and CD. InternVL2 follows closely, excelling in more complex tasks like AR, where it achieves a surprisingly 97.9% accuracy, far surpassing the RSVLMs. Additionally, InternVL2 shows strong reliability, reaching 80% accuracy on the HD task, well above others by a substantial margin.
RSVLMs outperform general-domain VLMs in advanced remote sensing tasks. For instance, the LHRS-Bot achieves a 36.0% accuracy in EA, and GeoRSCLIP reaches 43.0% in TP and 97% in DD tasks, surpassing all the general-domain VLMs. This indicates that RSVLMs hold a distinct advantage in understanding and analyzing complex RSIs.
@@ 4.3.2 Capability Dimensions
Current VLMs have yet to reach the ceiling for remote sensing capabilities. Even the highest-performing VLM, Qwen2-vl, achieves less than 60% accuracy on 40% of tasks in COREval, highlighting the substantial room for improvement in remote sensing capabilities among current VLMs.
Limited reasoning capabilities in remote sensing application scenarios. Critical reasoning tasks for remote sensing applications, such as PP, EA, and RA, remain challenging for most VLMs, with accuracies in these tasks generally falling below 30%. This suggests that current VLMs lack adequate development in reasoning abilities specifically tailored for remote sensing needs.
Visual grounding remains challenging. Visual grounding, a critical task for target localization in remote sensing, continues to pose significant difficulties, with existing RSVLMs achieving accuracy rates below 30%. General-domain VLMs perform even worse in this task, underscoring the need for further advancements.
Inadequate refusal capability. Except for InternVL2, the refusal capability among general-domain VLMs remains below 70%, warranting further enhancement. RSVLMs almost entirely lack refusal capabilities. Given the varying resolution and quality of RSIs in practical applications, the refusal for low-quality images and the ability to detect hallucinations deserves greater emphasis from researchers.
In summary, while general-domain VLMs have foundational strengths, effective applications in remote sensing require targeted field-specific enhancements. For RSVLMs, efforts should concentrate on achieving well-rounded improvements across all remote sensing tasks. COREval serves as a valuable resource for advancing the remote sensing capabilities of VLMs and guiding the refinement of these models’ specialized competencies.

### v2

@@ 4 Evaluation Results
4.1 Experimental Setup
We evaluated 24 mainstream VLMs based on CHOICE, which can be categorized into three groups: 15 open-source general-domain VLMs, including Qwen2-VL series [3], Ovis1.6-Gemma2-9B [40], InternVL2 series [5], LLaVA-1.6 series [41], Llama3.2 [4], GLM-4V [42], DeepSeek-VL [43], MiniCPM-V [44], Phi3-Vision [45], mPLUG-Owl3 [46] and Molmo [47], 6 RSVLMs: GeoChat [9], LHRS-Bot [10], LHRS-Bot-nova [48], VHM [24], RemoteCLIP [8], and GeoRSCLIP [23], and 3 proprietary VLMs: GPT-4o-mini, GPT-4o-2024-11-20 [49], and Gemini-1.5-Pro [50]. All VLMs in our evaluation employ the same zero-shot prompt strategy. Since most of our problems are MCQs, this format ensures the objectivity and rigor of the evaluation. For Visual Grounding task, we set IoU metric ($>$0.5 is considered correct) to compute the accuracy.
4.2 Main Results
The overall evaluation results of CHOICE are depicted in Figure 1 and Table 2. Currently, general-domain VLMs, benefiting from extensive training corpora, exhibit a broader range of capabilities, even outperforming RSVLMs in certain abilities pertinent to remote sensing. However, none of the VLMs achieve consistently high performance across the 6 capability dimensions of CHOICE, highlighting substantial room for improvement in tackling complex remote sensing tasks.
ILC and CSR Dimensions. The results for these two L-2 dimensions indicate that the majority of general-domain VLMs and RSVLMs exhibit strong capabilities in ILC and CSR, with task accuracy rates reaching approximately up to 80% or higher. This outcome demonstrates that these VLMs generally possess the foundational and critical capabilities for reasoning about common sense knowledge in remote sensing scenarios and comprehending RSIs at the image level.
SII and CID Dimensions. However, for the perception of finer-grained remote sensing targets, the results in Single-instance Identification and Cross-instance Discernment reveal that the accuracy of most general VLMs falls below 70%, while that of most RSVLMs is under 50%, indicating that current VLMs exhibit limitations in perception capabilities towards more complex remote sensing scenarios and local targets, necessitating further enhancements.
AttR and AssR Dimensions. Additionally, capabilities in AttR and AssR, which involve the analysis and reasoning of advanced attributes and features relevant to remote sensing, are largely underdeveloped in most existing VLMs. The evaluated models exhibit an accuracy of nearly 50%, with some RSVLMs even falling below 30%. However, these capabilities are of significant importance and value for the practical applications of large models in authentic remote sensing scenarios, underscoring the need for greater focus in this area.
Table 2: Evaluation results for L-2 dimensions. Abbreviations adopted: ILC for Image-level Comprehension; SII for Single-instance Identification; CID for Cross-instance Discernment; AttR for Attribute Reasoning; AssR for Assessment Reasoning; CSR for common sense Reasoning. The best (second best) is in bold (underline).
  Model ILC SII CID AttR AssR CSR
       Proprietary Large Vision-Language Models
GPT-4o-mini 0.800 0.588 0.448 0.494 0.474 0.876
GPT-4o-2024-11-20 0.845 0.616 0.591 0.536 0.277 0.900
Gemini-1.5-Pro 0.867 0.585 0.636 0.590 0.611 0.876
       General-domain Large Vision-Language Models
Qwen2-VL-7B 0.806 0.638 0.675 0.600 0.550 0.884
[…]
Ovis1.6-Gemma2-9B 0.828 0.606 0.632 0.598 0.645 0.890
[…]
LLaVA-1.6-7B 0.686 0.463 0.574 0.450 0.433 0.749
LLaVA-1.6-13B 0.719 0.522 0.626 0.474 0.470 0.733
Llama3.2-11B 0.746 0.502 0.504 0.460 0.388 0.803
GLM-4V-9B 0.785 0.555 0.644 0.564 0.450 0.863
DeepSeek-VL-7B 0.764 0.561 0.595 0.528 0.373 0.840
MiniCPM-V-2.5 0.757 0.546 0.631 0.414 0.493 0.857
Phi3-Vision 0.710 0.503 0.582 0.428 0.253 0.700
mPLUG-Owl3-7B 0.766 0.524 0.487 0.422 0.340 0.850
Molmo-7B-D 0.720 0.553 0.648 0.536 0.548 0.724
       Remote Sensing Large Vision-Language Models
GeoChat 0.642 0.469 0.480 0.384 0.368 0.697
LHRS-Bot 0.633 0.290 0.171 0.366 0.426 0.610
LHRS-Bot-nova 0.688 0.530 0.526 0.450 0.120 0.644
VHM 0.751 0.703 0.436 0.392 0.348 0.744
RemoteCLIP 0.657 0.283 0.552 0.326 0.364 0.739
GeoRSCLIP 0.745 0.198 0.285 0.210 0.397 0.884
  AttR AssR CSR
Model TP PP EA RA DD GD SI
       Proprietary Large Vision-Language Models
GPT-4o-mini 0.420 0.543 0.420 0.492 0.870 0.925 0.847
GPT-4o-2024-11-20 0.565 0.517 0.470 0.213 0.890 0.950 0.873
[…]
       General-domain Large Vision-Language Models
Qwen2-VL-7B 0.510 0.660 0.580 0.540 0.880 0.970 0.830
[…]
InternVL2-8B 0.520 0.760 0.400 0.640 0.880 0.790 0.790
InternVL2-26B 0.510 0.650 0.690 0.620 0.850 0.960 0.870
[…]
LLaVA-1.6-7B 0.300 0.550 0.320 0.470 0.740 0.710 0.780
LLaVA-1.6-13B 0.330 0.570 0.560 0.440 0.740 0.640 0.790
Llama3.2-11B 0.430 0.480 0.440 0.370 0.810 0.890 0.740
GLM-4V-9B 0.420 0.660 0.360 0.480 0.860 0.990 0.780
DeepSeek-VL-7B 0.390 0.620 0.170 0.440 0.860 0.880 0.800
[…]
Phi3-Vision 0.380 0.460 0.290 0.240 0.850 0.580 0.680
mPLUG-Owl3-7B 0.350 0.470 0.520 0.280 0.900 0.860 0.810
Molmo-7B-D 0.500 0.560 0.600 0.530 0.940 0.530 0.710
       Remote Sensing Large Vision-Language Models
GeoChat 0.255 0.470 0.105 0.455 0.660 0.810 0.647
LHRS-Bot 0.260 0.437 0.480 0.408 0.550 0.435 0.767
LHRS-Bot-nova 0.315 0.540 0.185 0.098 0.525 0.520 0.807
VHM 0.405 0.383 0.340 0.350 0.740 0.760 0.737
RemoteCLIP 0.335 0.210 0.035 0.245 0.820 0.650 0.740
[…]
4.3 Fine-grained Analysis
Table 4: Fine-grained evaluation results for Perception. Abbreviations adopted: IM for Image Modality; IQ for Image Quality; MR for Map Recognition; SC for Scene Classification; IC for Image Caption; LR for Landmark Recognition; OC for Object Counting; OL for Object Localization; OP for Object Presence; AR for Attribute Recognition; VG for Visual Grounding; HD for Hallucination Detection; AC for Attribute Comparison; SR for Spatial Relationship; CD for Change Detection.
@@ 4.3 Fine-grained Analysis
  Image-Level Comprehension Single-Instance Identification Cross-Instance Discerment
@@
       Proprietary Large Vision-Language Models
GPT-4o-mini 0.781 0.500 0.490 0.937 0.940 0.940 0.542 0.612 0.914 0.604 $0$ul0.906 0.426     0.340 0.650
[…]
@@
[…]
LLaVA-1.6-7B 0.340 0.460 0.380 0.920 0.880 0.740 0.220 0.690 0.970 0.590 0.237 0.060 0.730     0.340 0.650
LLaVA-1.6-13B 0.480 0.490 0.400 0.910 0.940 0.700 0.400 0.670 0.860 0.610 0.302 0.300 0.790     0.400 0.680
Llama3.2-11B 0.650 0.400 0.650 0.910 0.940 0.890 0.510 0.650 0.920 0.590 0.002 0.360 0.580     0.310 0.660
GLM-4V-9B 0.660 0.500 0.680 0.940 0.950 0.980 0.530 0.670 0.930 0.600 0.003 0.620 0.840     0.370 0.710
DeepSeek-VL-7B 0.570 0.540 0.670 0.920 0.940 0.890 0.460 0.740 0.950 0.530 0.253 0.430 0.770     0.340 0.670
MiniCPM-V-2.5 0.610 0.430 0.690 0.930 0.980 0.900 0.460 0.600 0.940 0.610 0.055 0.640 0.790     0.400 0.700
Phi3-Vision 0.510 0.480 0.480 0.880 0.910 0.660 0.350 0.760 0.910 0.560 0.105 0.380 0.770     0.370 0.570
mPLUG-Owl3-7B 0.770 0.420 0.670 0.890 0.960 0.920 0.500 0.660 0.950 0.590 0.073 0.380 0.520     0.300 0.710
Molmo-7B-D 0.560 0.530 0.350 0.870 0.920 0.740 0.390 0.630 0.920 0.670 0.015 0.760 0.800     0.490 0.620
@@
GeoChat 0.313 0.299 0.433 0.922 0.710 0.790 0.218 0.762 0.934 0.510 0.297 0.064 0.671     0.300 0.415
LHRS-Bot 0.288 0.306 0.318 0.915 0.755 0.500 0.252 0.164 0.936 0.330 - 0.076 -     0.267 0.325
[…]
VHM 0.621 0.428 0.299 0.966 0.765 0.760 0.342 0.872 0.932 0.564 0.598 0.922 0.431     0.347 0.580
RemoteCLIP 0.510 0.303 0.369 0.891 0.775 0.700 0.435 0.295 0.715 0.500 - 0.050 0.688     0.248 0.545
GeoRSCLIP 0.705 0.303 0.471 0.944 0.860 0.980 0.230 0.360 0.750 0.250 - 0.050 0.361     0.248 0.570
@@
Table 3 and 4 show the experimental results of general-domain VLMs and RSVLMs across all L-3 leaf tasks on CHOICE. See more experiments (blind and Circular evaluation [11]) in Appendix E.
@@ 4.3.1 Model Dimensions
General-domain VLMs match or even outperform RSVLMs in most L-3 tasks. The superior performance of VHM in SC, OL and VG demonstrates the efficacy of integrating domain-specific data to bridge the domain gap, but a clear overall disparity remains when compared to general-domain VLMs. Notably, Qwen2-VL-72B achieves the highest accuracy in a majority of perception tasks, exceeding 80% in MR, IC, OL, and CD. InternVL2-40B follows closely, excelling in more complex tasks like AC, TP and SI, far surpassing RSVLMs. Furthermore, RSVLMs remain confined to the perceptual level, trailing significantly behind general-domain VLMs in reasoning. For instance, while VHM showcases proficiency in Perception tasks, its reasoning capability is notably weaker than that of models with a similar parameter scale, like Ovis1.6-Gemma2-9B, GLM-4V-9B and Molmo-7B-D. We consistently observe that larger model capacities, exemplified by the Qwen2-VL and InterVL2 series, correlate with enhanced accuracy. These highlight the necessity of remote sensing data and that scaling parameters and incorporating general reasoning capability for RSVLMs can unlock additional performance gains.
Proprietary VLMs offer no clear advantage in remote sensing tasks. While three proprietary VLMs demonstrate top-tier capabilities in most L-3 tasks, open-source VLMs like Qwen2-VL-72B and InternVL2-40B show competitive or even superior performance. For example, Qwen2-VL-72B outperforms GPT-4o by 7% in CD and 15.3% in PP, respectively. These open-source VLMs emerge as attractive alternatives, offering a versatile and cost-effective solution for specific tasks.
@@ 4.3.2 Capability Dimensions
VLMs struggle with fine-grained perception. For tasks like OC, AR and SR, which require local fine-grained information and instance-level interactions, VLMs achieve an accuracy near or below 50%, clearly lower than other perception tasks. Visual grounding, a critical task for target localization in remote sensing, continues to pose significant difficulties, with existing RSVLMs achieving accuracy rates below 60%. General-domain VLMs perform even worse in this task, with some lacking this capability entirely. We attribute these shortcomings to the domain gap between RSIs and natural images, as well as the limitation of VLMs in extracting small-scale objects.
Reasoning capability is a bottleneck for VLMs. Critical reasoning tasks for remote sensing applications, such as TP, EA, and RA, remain challenging for most VLMs, with accuracies in these tasks generally falling below 50%, and even below 30% for some RSVLMs. This suggests that current VLMs lack adequate development in reasoning abilities tailored for remote sensing needs.
General-domain VLMs are more inclined to avoid model hallucination. In the HD task, we deliberately design problems that prompt VLMs to identify the locations of objects that do not exist in the images, thus testing their ability to mitigate model hallucination [51]. Except for VHM, which has been specifically trained on hallucination data, RSVLMs underperform most general-domain VLMs. However, the refusal capability among general-domain VLMs remains below 70%, which still falls short of the ideal threshold. Moreover, proprietary VLMs tend to favor the “won’t wrong” option when uncertain. For instance, GPT-4o selected “can’t judge” for 70.2% of the problems in the L-3 RA task, leading to a low accuracy of 0.277 in the L-2 AssR dimension, even worse than GPT-4o-mini. Given the varying resolution and quality of RSIs in practical applications, the refusal for low-quality images and the ability to detect hallucinations deserves greater emphasis from researchers.